Assume there is consumer, his utility function is u(x,y) =8*x^0.5+y , and his budget constraint is px*x +y = m, which implies py = 1.


a.Please derive the Marshallian demand function of x.

b.Please derive the indirect utility function.

c.Please derive the expenditure function.


1
Expert's answer
2021-10-28T08:49:37-0400

a) Marshallian demand function of x


Max "8x^{0.5} + y"


Subject to "Px^{x} + y = 1"


use budget constrain to determine y


"y = 1 - px^{x}"


Substituting in the utility function


max "f(x)= 8x^{0.5} + (1-Px^{x})"


Take first derivative using product rule


f(x) = "4x^{-0.5} + Px^{2}"


Set "f(x) = 0" and obtain


"4x^{-0.5} = - Px^{2}"


Solving for x


"x= \frac{-Px^{2.5} }{4}"



b) Indirect utility function

"U^*(Px,y,M) = max U(x,y)| px*x +y = m"


"=U(x^*,y^*)"


"= U(Dx(Px,Py,M), Dy(Px,Py,M))"



c) Expenditure function


"u = U^*(Px.Py,M) \iff M=M^*(Px,Py,u)"



"M^*(Px,Py,u) = min{Px^X+Py^y|U(x,y)}\ge u"

"x^* = D_x^H (Px,Py, u)" "y^* = D_y ^H (Px,Py,u)"